OpenStudy (icalibear):

Given the following sequence 2, 8, 18, 32, 50 ... what is the nth term in this sequence?

OpenStudy (icalibear):

a. 2n b. 2n^2 c. n + 1 d. n^2 + 1

OpenStudy (zzr0ck3r):

let n =1,2,3,4,... see which one works

OpenStudy (icalibear):

4?

OpenStudy (zzr0ck3r):

4 what?

OpenStudy (icalibear):

would n equal 4?

OpenStudy (zzr0ck3r):

take 2n^2 let n = 1, and you get 2*1=2 let n = 2 and you get 2*2^2 = 8 let n = 3 and you get 2*3^2 = 18 let n = 4.....

OpenStudy (icalibear):

2*4^2 = 64?

OpenStudy (bibby):

2*4^2 = 2*16 = 32

OpenStudy (icalibear):

Got it.